OpenStudy (anonymous):

2x-y=-4 -3x+y=-9 the answer is y=13 and x=30 show steps please

OpenStudy (jackellyn):

Solve for y in the first equation 2x -y = -4 -y = -4 -2x y= 4+2x Substitute into second equation -3x+(4+2x) = -9 Simplify/Combine like terms/Solve for x -3x +4 +2x =-9 -x +4 =-9 -x = -13 x=13 Substitute x=13 into first equation to find y 2(13)-y =-4 Simplify and solve for y 26 -y = -4 -y = -4-26 y =30

OpenStudy (jackellyn):

I got opposite answers.

OpenStudy (anonymous):

i probully did it wrong how would you do a check step for this?

OpenStudy (jackellyn):

if you substitute the value you got for x and y in either equation they should equal the number it says. for example: 2(13)-30=-4 26-30 =-4 -4=-4

OpenStudy (anonymous):

sweeet thank you

OpenStudy (jackellyn):

no problem :)